# Beyond Position: the emergence of wavelet-like properties in Transformers

### v3

@@ Abstract
This paper studies how transformer models develop robust wavelet-like properties that effectively compensate for the theoretical limitations of Rotary Position Embeddings (RoPE), providing insights into how these networks process sequential information across different scales. Through theoretical analysis and empirical validation across models ranging from 1B to 12B parameters, we show that attention heads naturally evolve to implement multi-resolution processing analogous to wavelet transforms. Our analysis establishes that attention heads consistently organize into complementary frequency bands with systematic power distribution patterns, and these wavelet-like characteristics become more pronounced in larger models. We provide mathematical analysis showing how these properties align with optimal solutions to the fundamental uncertainty principle between positional precision and frequency resolution. Our findings suggest that the effectiveness of modern transformer architectures stems significantly from their development of optimal multi-resolution decompositions that naturally address the theoretical constraints of position encoding.
@@
Valeria Ruscio Sapienza University of Rome ruscio@diag.uniroma1.it                        Fabrizio Silvestri Sapienza University of Rome fsilvestri@diag.uniroma1.it
@@ 1 Introduction
Position encoding mechanisms are fundamental to transformer architectures, enabling these inherently permutation-invariant models to capture sequential information crucial for natural language understanding. While early approaches relied on fixed sinusoidal encodings (Vaswani, 2017), Rotary Positional Embeddings (RoPE) (Su et al., 2024) represents a significant advancement by directly integrating positional information through learned rotations of token embeddings. RoPE’s elegant mathematical properties and demonstrated effectiveness have led to its widespread adoption in state-of-the-art language models.
Despite RoPE’s success in practice, theoretical analysis suggests certain inherent limitations. Like all position encoding schemes, RoPE faces fundamental trade-offs between positional precision and frequency resolution, analogous to the uncertainty principle in signal processing. However, what makes RoPE particularly interesting is how these theoretical limitations seem to have little practical impact on model performance, as seen in Barbero et al. (2024). This discrepancy between theoretical constraints and empirical success motivates our investigation into how transformer models adapt to and overcome these apparent limitations.
Our analysis reveals that transformer attention heads, develop sophisticated wavelet-like properties that effectively address these theoretical constraints. Different heads naturally specialize in processing information at distinct frequency bands, creating a multi-resolution framework that balances local and global information processing. This organization emerges consistently across model scales, suggesting it represents a fundamental property of how neural networks optimize position-aware sequence processing.
Through mathematical analysis and extensive empirical validation, we establish several key connections between RoPE-based attention mechanisms and wavelet transforms. The attention patterns that emerge during training show remarkable similarity to wavelet basis functions, with heads automatically organizing into complementary frequency bands. This specialization provides an adaptive decomposition of input sequences that elegantly balances the theoretical trade-offs inherent in position encoding.
Our work makes two main contributions:
  • We provide a theoretical framework connecting RoPE-based attention mechanisms with wavelet theory, offering new insights into how transformers process sequential information.
  • We demonstrate through empirical analysis how attention heads develop wavelet-like properties that effectively address theoretical limitations.
These findings reveal that transformers naturally evolve sophisticated mechanisms for multi-resolution analysis of sequential data. Rather than highlighting limitations, our work underscores the remarkable adaptability of neural architectures in developing optimal solutions to complex information processing challenges. This understanding opens new avenues for architectural innovation while deepening our appreciation of existing approaches.
@@ 2 Related Works
The Transformer architecture Vaswani (2017) revolutionized sequence modeling by introducing self-attention mechanisms, eliminating the need for recurrent structures. The original Transformer used additive sinusoidal positional encodings, which provided a simple but effective way to inject position information.
Recent work has explored more sophisticated approaches to position encoding. ALiBi (Press et al., 2021) introduced attention bias terms that scale with relative position, while T5 (Raffel et al., 2020) employed learned relative position embeddings. Rotary Position Embedding (RoPE) (Su et al., 2024) represented a significant advancement by applying rotation matrices to embeddings, introducing relative positional dependence through phase shifts while preserving inner product structures.
RoPE encodes positional information by applying rotation matrices to token embeddings in the complex plane, where the rotation angle is a function of position and frequency. While this approach elegantly preserves the inner product between tokens while encoding their relative positions, it faces a fundamental limitation rooted in the uncertainty principle: it cannot simultaneously achieve perfect precision in both position and frequency domains. Theoretically, this suggests RoPE should struggle with tasks requiring both precise positional information and broad frequency understanding. However, in practice, transformer models achieve remarkable performance despite this theoretical constraint.
The behavior of neural networks, particularly their nonlinear components, has been increasingly analyzed through the lens of signal processing. Research has shown that activation functions like ReLU and GeLU can generate higher-order harmonics and exhibit frequency mixing (Selesnick and Burrus, 1998; Rahimi and Recht, 2008). These effects become particularly relevant in understanding how positional information propagates through transformer networks.
Principles of constructive and destructive interference from signal processing (Oppenheim, 1999) have proven valuable in analyzing neural network behavior. Recent work has examined how neural networks process frequency information (Chi et al., 2020), while others have drawn parallels between neural activations and signal modulation techniques (Takahashi et al., 2018; Mildenhall et al., 2021).
Information-theoretic analyses of neural networks (Shwartz-Ziv and Tishby, 2017) have provided insights into their representational capabilities and limitations. Studies have examined how information flows through layers (Goldfeld et al., 2018) and how architectural choices affect information bottlenecks (Tishby and Zaslavsky, 2015). This theoretical framework has proven particularly valuable in understanding the capacity limitations of various neural network components.
@@ 3 Methodology
In this section, we describe the methodological framework employed to investigate how Transformer models utilizing Rotary Position Embeddings (RoPE) develop compensatory mechanisms that transcend their theoretical positional encoding limitations. We integrate frequency-domain analyses, wavelet-based multi-scale decomposition, and entropy-based uncertainty assessments to comprehensively characterize the emergent properties of these models. Our methodology is designed to isolate positional encoding behaviors, assess their stability across model scales and architectures, and validate their alignment with theoretical expectations related to the trade-off between positional resolution and spectral organization.
@@ 3.1 Frequency Analysis
To probe the spectral properties of attention distributions, we used a frequency-domain analysis using the Discrete Fourier Transform (DFT). For each attention head $h$ within each model, we represented the attention pattern over token positions as $a_{h}(t)$, where $t$ indexes tokens within a single sequence. We computed the power spectral density (PSD):
@@
where $b$ is the frequency band under consideration. These frequency-domain analyses allowed us to discern how attention heads distribute their representational capacity across multiple scales, testing the premise that models spontaneously develop organized frequency content despite RoPE’s intrinsic limitations.
@@ 3.2 Wavelet Analysis
While frequency-domain analysis captures global spectral properties, it lacks explicit positional localization. To address this, we used wavelet decompositions using the Daubechies-2 (db2) wavelet. Wavelets offer a time-frequency (or position-frequency) representation that enables simultaneous assessment of spatial localization and scale-dependent behaviors.
@@ 3.3 Uncertainty Analysis
By comparing $H_{p}(h)$ and $H_{s}(h)$, we can ascertain whether the model’s attention patterns obey an uncertainty principle-like trade-off, wherein improved positional localization may come at the cost of reduced spectral complexity, or vice versa.
@@ 3.4 Scale Invariance Testing
We hypothesized that the models’ compensatory strategies would exhibit scale invariance properties, that is, the ability to maintain positional awareness structures when the input sequence length changes. To test this, we generated scaled variants $x_{\alpha}$ of each input sequence $x$ by sampling $\lfloor\alpha n\rfloor$ tokens, with $\alpha\in\{0.5,0.25\}$ and $n$ the original sequence length. After computing the wavelet coefficients $W_{h}(x)$ and $W_{h}(x_{\alpha})$, we measured the scale sensitivity:
@@ 3.5 Frame Completeness
where $W^{-1}(\cdot)$ denotes the inverse wavelet transform and $||\cdot||_{F}$ is the Frobenius norm. A small $\varepsilon$ indicates that the attention patterns are well represented by their wavelet coefficients, reinforcing the notion that the model’s positional strategies form a coherent, frame-like structure.
@@ 4 Implementation Details
We selected five pre-trained Transformer-based language models that vary in size, architecture, and training regimen to ensure the generality of our findings. Specifically, we analyzed Gemma 2 2B, Pythia 2.8B and 12B, LLaMA-3-2 1B, Mistral 7B, and Qwen 2.5 5B. These models cover a wide parameter range (1B–12B), capturing different representational capacities and training protocols.
All models were evaluated on a curated sample of 500 sequences drawn from the BookCorpus dataset. Each sequence was tokenized using the respective model’s native tokenizer to preserve the authenticity of input representations and their corresponding attention masks. The selected sequences varied in length to expose scale-dependent behavior and stress-test the models’ positional encoding strategies under diverse conditions.
All experiments were performed using PyTorch on A100, L4 and T4 GPUs to ensure computational efficiency and scalability. Frequency and spectral computations employed standard FFT-based routines, while wavelet transforms were performed using the PyWavelets library with a decomposition level chosen based on the minimum sequence length. Before analysis, attention weights were normalized and numerically stabilized to mitigate floating-point underflow, with a threshold of $10^{-10}$ applied to division operations.
@@ 5 Experiments and Analysis
Our empirical analysis reveals striking patterns in how transformer models organize their attention mechanisms to process information across different scales.
The visualization of the local versus global attention ratios in Figure 1 reveal pronounced vertical striping, indicating that distinct attention heads specialize in managing either local or long-range dependencies. Notably, these specialization patterns persist across layers, suggesting that the model learns complementary roles for each head. Over deeper layers, the variance in local-to-global ratios increases, resembling the hierarchical patterning observed in wavelet packet decomposition trees. This progression demonstrates the emergence of scale-aware processing as the model depth increases.
@@
Our frequency band distribution visualizations in Figure 2 highlight a hierarchical structure in how attention heads allocate their representational capacity across spectral components. The low-frequency range (0–0.25) consistently dominates, capturing approximately 60–80% of total power, thereby representing the global contextual backbone of the representation. Mid-frequency components (0.25–0.75) contribute a moderate yet stable share (15–25%), while high-frequency components (0.75–1.0) maintain a smaller but non-negligible presence (5–15%). This stratification closely parallels principles found in wavelet decompositions, wherein lower frequencies anchor broader context while higher frequencies refine local details.
@@
The temporal evolution of frequency responses in Figure 3 gives us further evidence for wavelet-like properties. at the beginning, low-frequency dominance gradually tapers, while mid- and high-frequency components gain influence. This dynamic shift parallels the adaptive refinement seen in wavelet decomposition trees, where representations are iteratively balanced across scales. Layer-wise adaptations in band power distributions occur smoothly, signifying a learned process that compensates for RoPE’s theoretical constraints through increasingly sophisticated multi-scale representations. Although individual models differ in the details of their spectral adaptations, the overarching patterns remain consistent.
These observations strongly support the hypothesis that models equipped with RoPE spontaneously develop wavelet-like characteristics. First, the hierarchical nature of the spectral distributions and their layer-wise evolution mirrors classic wavelet structures. Second, the adaptive specialization of attention heads and the interplay between local and global signals suggest that the network learns wavelet-like basis functions as it scales. Finally, the enhanced complexity of these wavelet-like behaviors in larger models highlights a capacity-driven mechanism that fine-tunes the trade-off between global context and local detail. Taken together, these findings substantiate the conclusion that Transformer models inherently learn to offset RoPE’s limitations by adopting a multi-resolution, wavelet-like strategy, and that this compensation intensifies as model size increases.
@@
Model Heads Spectral Frequency Scale 0.5 Scale 0.25 Pos-Spec Reconstr.

### v4

@@ Abstract
This paper studies how Transformer models with Rotary Position Embeddings (RoPE) develop emergent, wavelet-like properties that compensate for the positional encoding’s theoretical limitations. Through an analysis spanning model scales, architectures, and training checkpoints, we show that attention heads evolve to implement multi-resolution processing analogous to wavelet transforms. We demonstrate that this scale-invariant behavior is unique to RoPE, emerges through distinct evolutionary phases during training, and statistically adheres to the fundamental uncertainty principle. Our findings suggest that the effectiveness of modern Transformers stems from their remarkable ability to spontaneously develop optimal, multi-resolution decompositions to address inherent architectural constraints.
@@
Valeria Ruscio, Umberto Nanni, Fabrizio Silvestri Sapienza University of Rome ruscio@diag.uniroma1.it, fsilvestri@diag.uniroma1.it
@@ 1 Introduction
Position encoding mechanisms are fundamental to Transformer architectures, enabling these inherently permutation-invariant models to capture sequential information. While early approaches relied on fixed sinusoidal encodings (Vaswani, 2017), Rotary Positional Embeddings (RoPE) (Su et al., 2024) represents a significant advancement by integrating relative positional information through rotational transformations. Despite RoPE’s widespread adoption and empirical success, theoretical analysis suggests inherent limitations in its ability to simultaneously achieve high positional precision and frequency resolution (Barbero et al., 2024), a trade-off analogous to the uncertainty principle in signal processing. This creates a paradox: why do models with these theoretical constraints perform so well in practice?
We argue that this is resolved because RoPE-equipped models learn to compensate by developing emergent, wavelet-like processing strategies. Our analysis shows that Transformer attention heads do not act as a monolithic block; instead, they spontaneously organize into a multi-resolution framework. Different heads specialize in processing information at distinct frequency bands, effectively decomposing the input signal in a manner strikingly similar to a discrete wavelet transform.
Our work makes the following contributions:
  • We demonstrate that the emergence of robust, wavelet-like, scale-invariant properties is a distinctive feature of the RoPE architecture compared to other common position encoding schemes.
  • We provide an evolutionary analysis of these properties, revealing the distinct learning phases models undergo to form their spectral processing strategies.
  • We offer a detailed statistical characterization of attention head behavior, showing that models learn to respect the physical constraints of the uncertainty principle and develop diverse internal strategies.
These findings reveal transformers’ remarkable adaptability in discovering and implementing optimal solutions to complex information processing challenges. Appendices 11.3, 11.4, and 11.5 provide further detail on RoPE’s limitations, the relationship between language and wavelets, and our metric definitions.
@@ 2 Related Works
The Transformer architecture Vaswani (2017) revolutionized sequence modeling through self-attention mechanisms. While the original Transformer used simple sinusoidal positional encodings, recent work has explored more sophisticated approaches. ALiBi (Press et al., 2021) introduced attention bias terms that scale with relative position, while T5 (Raffel et al., 2020) employed learned relative position embeddings. RoPE (Su et al., 2024) advanced this further by applying rotation matrices to embeddings, though it faces fundamental limitations rooted in the uncertainty principle between position and frequency domains.
Neural networks’ behavior, particularly their nonlinear components, has been increasingly analyzed through signal processing principles. Research has shown that activation functions can generate higher-order harmonics and exhibit frequency mixing (Selesnick and Burrus, 1998; Rahimi and Recht, 2008), while principles of constructive and destructive interference have proven valuable in analyzing network behavior (Oppenheim, 1999; Chi et al., 2020). Information-theoretic analyses of neural networks (Shwartz-Ziv and Tishby, 2017) have provided insights into their representational capabilities and limitations. Studies have examined how information flows through layers (Goldfeld et al., 2018) and how architectural choices affect information bottlenecks (Tishby and Zaslavsky, 2015). This theoretical framework has proven particularly valuable in understanding the capacity limitations of various neural network components.
@@ 3 Methodology
We integrate frequency-domain analyses, wavelet-based multi-scale decomposition, and entropy-based uncertainty assessments to comprehensively characterize the emergent properties of these models. Our methodology is designed to isolate positional encoding behaviors, assess their stability across model scales and architectures, and validate their alignment with theoretical expectations related to the trade-off between positional resolution and spectral organization.
@@ 3.1 Frequency Analysis
To probe the spectral properties of attention distributions, we employed a frequency-domain analysis using the Discrete Fourier Transform (DFT), we used the Hann window and zero padding. For each attention head $h$ within each model, we represented the attention pattern over token positions as $a_{h}(t)$, where $t$ indexes tokens within a single sequence. We computed the power spectral density (PSD):
@@
The Nyquist frequency $\omega_{N}$ is set to half the sampling rate (1/2 tokens) for three fundamental reasons: it represents the highest meaningful frequency in discrete token sequences, as attention patterns can only alternate between consecutive tokens, making faster oscillations indistinguishable due to aliasing. Second, it provides natural normalization across sequence lengths, while absolute frequency ranges differ, all sequences share the same relative frequency structure when normalized by $\omega_{N}$, enabling meaningful cross-length comparisons of attention head frequency sensitivity. Third, following Shannon’s sampling theorem, $\omega_{N}$ represents the theoretical maximum rate for information transmission through a discrete channel, thus defining the finest granularity at which positional information can be encoded without loss, making it the natural choice for analyzing models’ representational capacity distribution.
@@
where $b$ is the frequency band under consideration. To measure how selectively each attention head responds to specific frequencies, we define the frequency selectivity $S(h)$ for head $h$ as:
[…]
where $P_{h}(\omega)$ is the power spectral density at frequency $\omega$, and $\omega N$ is the Nyquist frequency, and a higher value indicates more focused frequency tuning of the head.
These frequency-domain analyses allowed us to discern how attention heads distribute their representational capacity across multiple scales, testing the premise that models spontaneously develop organized frequency content despite RoPE’s intrinsic limitations.
@@ 3.2 Wavelet Analysis
While frequency-domain analysis captures global spectral properties, it lacks explicit positional localization. To address this, we employed wavelet decompositions using the Daubechies-2 (db2) wavelet  [^1] The selection of Daubechies-2 (db2) for our primary analysis reflects its optimal balance between smoothness and localization properties. Its compact support of length 4 aligns well with typical attention spans in language processing, while its vanishing moment enables effective detection of local linguistic changes against broader contextual background. The db2 wavelet’s mathematical normalization properties match the normalization constraints imposed by softmax attention mechanisms, while its orthogonality properties prevent interference between shifted attention patterns.. Wavelets offer a time-frequency (or position-frequency) representation that enables simultaneous assessment of spatial localization and scale-dependent behaviors.
@@ 3.3 Uncertainty Analysis
where $\hat{P}_{h}(\omega)=\frac{P_{h}(\omega)}{\sum_{\omega}P_{h}(\omega)}$ is the normalized power spectrum.
To quantify the relationship between these entropy measures, we define the position-spectrum correlation $\rho(h)$ through their normalized covariance:
$\rho(h)=\frac{\text{Cov}(H_{p}(h),H_{s}(h))}{\sigma_{H_{p}}\sigma_{H_{s}}}$ (8)
This correlation is then aggregated across all attention heads in a layer to measure how well the model balances the uncertainty principle trade-off between positional and spectral information:
$\rho_{\text{layer}}=\text{mean}_{h\in\text{layer}}\{\rho(h)\}$ (9)
The layer-wise correlation metric is bounded by $[-1,1]$, with values closer to -1 indicating strong trade-offs between positional and spectral precision, and values closer to 1 indicating successful integration of both domains.
By comparing $H_{p}(h)$ and $H_{s}(h)$ through these correlation metrics, we can ascertain whether the model’s attention patterns obey an uncertainty principle-like trade-off, wherein improved positional localization may come at the cost of reduced spectral complexity, or vice versa.
@@ 3.4 Scale Invariance Testing
We hypothesized that the models’ compensatory strategies would exhibit scale invariance properties, i.e., the ability to maintain positional-awareness structures when the input sequence length changes. To test this, we generated scaled variants $x_{\alpha}$ of each input sequence $x$ by sampling $\lfloor\alpha n\rfloor$ tokens, with $\alpha\in\{0.5,0.25\}$ and $n$ the original sequence length. After computing the wavelet coefficients $W_{h}(x)$ and $W_{h}(x_{\alpha})$, we measured the scale sensitivity:
@@ 3.5 Frame Completeness
where $W^{-1}(\cdot)$ denotes the inverse wavelet transform and $||\cdot||_{F}$ is the Frobenius norm. A small $\varepsilon$ indicates that the attention patterns are well-represented by their wavelet coefficients, reinforcing the notion that the model’s positional strategies form a coherent, frame-like structure.
@@ 4 Implementation Details
We selected five pre-trained Transformer-based language models that vary in size, architecture, and training regimen to ensure the generality of our findings. Specifically, we analyzed Gemma 2 2B, Pythia 2.8B and 12B, LLaMA-3-2 1B, Mistral 7B, and Qwen 2.5 5B. These models encompass a wide parameter range (1B–12B), capturing different representational capacities and training protocols.
All models were evaluated on a curated sample of 500 sequences drawn from wikipedia. The selected sequences varied in length to expose scale-dependent behavior and stress-test the models’ positional encoding strategies under diverse conditions.
All experiments were conducted using PyTorch on A100, L4, and T4 GPUs to ensure computational efficiency and scalability. Frequency and spectral computations employed standard FFT-based routines, while wavelet transforms were performed using the PyWavelets library with a decomposition level chosen based on the minimum sequence length. Before analysis, attention weights were normalized and numerically stabilized to mitigate floating-point underflow, with a threshold of $10^{-10}$ applied to division operations.
@@ 5 Experiments and Analysis
Our analysis shows that Transformer models with RoPE spontaneously develop a sophisticated, multi-resolution processing strategy, similar to wavelet decomposition, to overcome the theoretical limitations of their position embeddings. This emergent behavior is not an isolated artifact but a consistent pattern substantiated by three key lines of evidence: the hierarchical organization of attention, quantitative analysis of model scaling and stability, and statistical confirmation of a fundamental signal processing principle.
The Emergence of Multi-Scale Processing
In figure 1 we can see that attention heads specialize into either local or global processors, evidenced by the pronounced vertical striping in visualizations of local-to-global attention ratios. This specialization deepens through the layers, with increasing variance that closely resembles the branching structure of a wavelet packet decomposition tree.
This multi-resolution strategy is also evident in the frequency domain. Analysis of attention patterns, as in figure 2, shows a consistently stratified frequency response. Low-frequency components (0–0.25 $\omega N$) form the contextual backbone, capturing 60–80% of the spectral power. Mid-frequencies (0.25–0.75 $\omega N$) contribute a stable 15–25%, while high-frequencies (0.75–$\omega N$) handle fine-grained details with a smaller 5–15% share. In figure 3, we can see that as information propagates through the model, we observe a smooth, dynamic evolution where the initial dominance of low frequencies tapers off, and mid- to high-frequency components gain influence, mirroring the adaptive refinement process of a wavelet analysis.
@@
5.1 Quantitative Analysis of Developmental Trajectories
Quantitative metrics confirm these visual observations and reveal deeper patterns related to model scale and capacity, as we can see from table 1. The model’s inherent wavelet-like response is particularly evident in the scale sensitivity metric. As predicted by wavelet theory, sensitivity to sequence rescaling is not only low but also degrades gracefully, with the error roughly doubling as scaling moves from 0.5x to 0.25x. This predictable, controlled divergence mirrors how high-quality wavelet basis functions respond to dilation and suggests a learned adherence to power-law scaling behavior.
Furthermore, we observe a fascinating, non-monotonic developmental trajectory as model size increases. Mid-sized models like Pythia 2.8B appear to enter a unique "transitional exploration phase," marked by remarkably low frequency selectivity (0.174) and the highest spectral entropy (3.786). This suggests a strategy based on highly distributed, spectrally diffuse representations. In contrast, both smaller models (like Qwen 0.5B) and larger, more advanced models (like Mistral 7B) converge on higher selectivity, indicating they find a more optimized and efficient balance between specialized frequency channels and integrated multi-resolution analysis, a hallmark of mature wavelet-like systems.
This multi-resolution coherence is also confirmed by our window entropy analysis, shown in table 2. Most models demonstrate remarkable stability, indicating their representations remain coherent across different observational scales. Pythia 2.8B is again a notable exception, with its entropy decreasing as the window size grows $(0.751\rightarrow 0.624)$. This, combined with its other unique metrics, suggests a distinct hierarchical strategy where information is more complex and dense in local contexts but more predictable at broader scales.
@@ 5.1 Quantitative Analysis of Developmental Trajectories
Model Heads Scale Sens. (0.5x) Scale Sens. (0.25x) Pos-Spec Corr. ($\rho$) Reconstr. Error ($\mu$)
Qwen 2.5 (0.5B) 14 0.058 0.100 -0.438 1.26e-07
LLaMA 3.2 (1B) 32 0.038 0.089 -0.510 1.28e-07
Gemma 2 (2B) 8 0.038 0.090 -0.526 1.27e-07
Pythia (2.8B) 32 0.082 0.121 -0.737 1.16e-07
Mistral (7B) 32 0.030 0.074 -0.421 1.41e-07
LLaMA 3.1 (8B) 32 0.038 0.090 -0.474 1.28e-07
Pythia (12B) 40 0.059 0.099 -0.490 1.26e-07
Table 1: This table summarizes the main findings for scale sensitivity, position-spectrum correlation, and reconstruction error.
Model 16 tok. 32 tok. 64 tok.
LLaMA 3.2 (1B) 0.877 0.886 0.882
Gemma 2 (2B) 0.894 0.905 0.904
Pythia (2.8B) 0.751 0.699 0.624
Qwen 2.5 (0.5B) 0.850 0.864 0.866
Mistral (7B) 0.870 0.878 0.880
LLaMA 3.1 (8B) 0.869 0.877 0.871
Pythia (12B) 0.913 0.918 0.916
Table 2: Multi-Resolution Window Entropy Analysis
5.2 Statistical Confirmation of the Uncertainty Principle
The most direct evidence for this emergent behavior comes from a statistical analysis of the trade-off between positional and spectral information. As shown in table 3, every model analyzed exhibits a consistently negative position-spectrum correlation $\rho$. This finding is a powerful confirmation that the models have implicitly learned and adhere to the Heisenberg-Gabor uncertainty principle, a fundamental law governing all time-frequency representations, including wavelets.
This distributional analysis uncovers two distinct strategic archetypes. Models like Mistral 7B exhibit a high-variance, specialized toolkit approach. It pairs high average frequency selectivity ($\mu=0.804$) with a very wide standard deviation ($\sigma=0.414$), indicating that its attention heads are highly diversified for a sophisticated division of labor. In stark contrast, Pythia 2.8B adopts a low-variance, uniform strategy. Its extremely low mean frequency selectivity ($\mu=0.174$) is highly consistent across its heads (IQR=0.153), confirming its "transitional phase" is a model-wide phenomenon.
The varying strength of the position-spectrum correlation, from Pythia 2.8B’s severe trade-off ($\rho=-0.737$) to Mistral’s more efficient balance ($\rho=-0.421$), shows how different architectures have converged with varying degrees of success on the optimal, wavelet-like strategies for balancing the fundamental "what" versus "where" information trade-off inherent in sequence processing.
Model Spectral Entropy Frequency Selectivity
Mean ($\mu$) Std Dev ($\sigma$) IQR (Q3-Q1) Mean ($\mu$) Std Dev ($\sigma$) IQR (Q3-Q1)
Qwen 2.5 (0.5B) 2.492 0.733 0.985 0.700 0.411 0.498
LLaMA 3.2 (1B) 2.370 0.650 0.924 0.758 0.368 0.494
Gemma 2 (2B) 2.244 0.599 0.781 0.794 0.344 0.430
Pythia (2.8B) 3.786 0.625 0.398 0.174 0.269 0.153
Mistral (7B) 2.443 0.607 0.824 0.804 0.414 0.515
LLaMA 3.1 (8B) 2.495 0.621 0.859 0.691 0.339 0.444
Pythia (12B) 2.783 0.638 0.850 0.552 0.321 0.391
Table 3: Distributional statistics for spectral metrics across all attention heads. The standard deviation ($\sigma$) and interquartile range (IQR) reveal the diversity of learned strategies within each model.
5.3 Ablation Study
Model PE Type Heads Spectral Frequency Scale 0.5 Scale 0.25 Pos-Spec Reconstr.